OpenStudy (anonymous):

If the midpoint between (x, 6) and (-9, 14) is (8, 10), find the value of x.

OpenStudy (anonymous):

(-9+x)/2=8 -9+x=16 x=25

OpenStudy (anonymous):

Thaks

OpenStudy (anonymous):

Sorry....I mean Thanks